OpenStudy (anonymous):

what is the value of csc(-19π/3)? I have no trig. knowledge except for sohcahtoa

OpenStudy (kc_kennylau):

Do you know 1/csc is which trigo function?

OpenStudy (anonymous):

….not really but is it cos?

OpenStudy (kc_kennylau):

The first step is like this: \(\Large\displaystyle\csc\left(\frac{-19\pi}3\right)=\frac1{\sin\left(\frac{-19\pi}3\right)}\)

OpenStudy (kc_kennylau):

Remember that cosine is to secant and sine is to cosecant

OpenStudy (anonymous):

ok…. question, what is secant and cosecant… sorry

OpenStudy (kc_kennylau):

secant is 1/cosine, abbreviated as sec

OpenStudy (kc_kennylau):

cosecant is 1/sine, abbreviated as csc or cosec

OpenStudy (kc_kennylau):

Get it?

OpenStudy (anonymous):

i think so Thanks! :D so after csc(−19π3)=1sin(−19π3) what happens?

OpenStudy (kc_kennylau):

Do you know what's the period of sine?

OpenStudy (anonymous):

no

OpenStudy (kc_kennylau):

Okay, so for any trigonometric function, the period is always 2π, meaning that the function will repeat itself after 2π.

OpenStudy (kc_kennylau):

That means that \(\sin x=\sin(2\pi+x)\)

OpenStudy (kc_kennylau):

Get it?

OpenStudy (anonymous):

i think so

OpenStudy (kc_kennylau):

Okay, use the formula above 3 times

OpenStudy (anonymous):

it will get easier as i see it more but your doing an amazing job explaining to me and answering by never ending questions… ok

OpenStudy (kc_kennylau):

No problem, asking questions is the most effective way of learning

OpenStudy (anonymous):

so is it set up like this: csc(−19π3)=1sin(−19π3) X sine(2pi + x) X sine(2pi +x) X sin(2pi +x) ?

OpenStudy (kc_kennylau):

Never mind...

OpenStudy (kc_kennylau):

\[\Large\displaystyle\csc\left(\frac{-19\pi}3\right)=\frac1{\sin\left(\frac{-19\pi}3\right)}=\frac1{\sin\left(\frac{-19\pi}3+2\pi+2\pi+2\pi\right)}\]

OpenStudy (kc_kennylau):

\[\Large=\frac1{\sin\frac{-\pi}3}\]

OpenStudy (anonymous):

oh… ok thanks :D

OpenStudy (kc_kennylau):

You have the table right

OpenStudy (anonymous):

is that all there is to it?

OpenStudy (kc_kennylau):

Well you have to find its value using the table

OpenStudy (anonymous):

ok thanks you have been a great help!

OpenStudy (kc_kennylau):

No problem :)